# LinearRAG: Linear Graph Retrieval Augmented Generation on Large-scale Corpora

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) is widely used to mitigate hallucinations of Large Language Models (LLMs) by leveraging external knowledge. While effective for simple queries, traditional RAG systems struggle with large-scale, unstructured corpora where information is fragmented. Recent advances incorporate knowledge graphs to capture relational structures, enabling more comprehensive retrieval for complex, multi-hop reasoning tasks. However, existing graph-based RAG (GraphRAG) methods rely on unstable and costly relation extraction for graph construction, often producing noisy graphs with incorrect or inconsistent relations that degrade retrieval quality. In this paper, we revisit the pipeline of existing GraphRAG systems and propose Linear Graph-based Retrieval-Augmented Generation (LinearRAG), an efficient framework that enables reliable graph construction and precise passage retrieval. Specifically, LinearRAG constructs a relation-free hierarchical graph, termed Tri-Graph, using only lightweight entity extraction and semantic linking, avoiding unstable relation modeling. This new paradigm of graph construction scales linearly with corpus size and incurs no extra token consumption, providing an economical and reliable indexing of the original passages. For retrieval, LinearRAG adopts a two-stage strategy: (i) relevant entity activation via local semantic bridging, followed by (ii) passage retrieval through global importance aggregation. Extensive experiments on four datasets demonstrate that LinearRAG significantly outperforms baseline models.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a promising approach to enhance Large Language Models (LLMs) by leveraging external knowledge bases (Gao et al., 2023; Lewis et al., 2020; Zhou et al., 2025a; Zhang et al., 2025c). However, existing RAG systems struggle with the complexities of large-scale, unstructured corpora in real-world scenarios, where the relevant information is frequently distributed unevenly across heterogeneous documents. The context retrieved by RAG systems is often voluminous, intricate, and lacks clear organization, leading to issues of variability in accuracy and coherence  (Sun et al., 2024; Zhang et al., 2024). Although recent advances attempt to manage this by segmenting documents into smaller chunks for efficient indexing (Borgeaud et al., 2022; Izacard et al., 2023; Jiang et al., 2023), this strategy often results in the loss of critical contextual details, impairing retrieval accuracy and reasoning capabilities for complex tasks (Han et al., 2024; Zhang et al., 2025b).
To address this, Graph Retrieval-Augmented Generation (GraphRAG) (Zhang et al., 2025b; Procko & Ochoa, 2024; Xiao et al., 2025a) has recently emerged as a powerful paradigm that leverages external structured graphs to model the hierarchical structure of background knowledge (Han et al., 2024). Specifically, early work, like RAPTOR (Sarthi et al., 2024) and Microsoft’s GraphRAG (Edge et al., 2024), organizes knowledge through recursive summarization and community detection with LLM-generated synopses, enabling coarse-to-fine retrieval for comprehensive responses. Building on this, recent approaches, including GFM-RAG (Luo et al., 2025), G-Retriever (He et al., 2024), and LightRAG (Guo et al., 2024), integrate specialized encoders and objectives, such as query-dependent GNNs, Prize Collecting Steiner Trees, and dual-level indexing, to improve multi-hop generalization, scalability, and efficiency. More recently, HippoRAG (Gutiérrez et al., 2024) and its enhancement HippoRAG2 (Gutiérrez et al., 2025) draw inspiration from cognitive processes to utilize personalized PageRank for multi-hop retrieval. These strategies significantly improve retrieval precision and contextual depth, enabling LLMs to address complex, multi-hop queries more effectively.
@@
Despite its conceptual promise and theoretical superiority, recent studies reveal that GraphRAG models frequently underperform even naive RAG approaches on many real-world applications (Han et al., 2025; Zhou et al., 2025b; Xiang et al., 2025; Zhuang et al., 2025). This performance degradation mainly stems from the poor quality of automatically constructed knowledge graphs. While graph-based retrieval increases recall of relevant knowledge, it concurrently introduces substantial noise and ambiguities into the retrieved contexts, due to errors in graph construction. Specifically, two critical deficiencies undermine graph quality, including (i) local inaccuracy: relation extraction processes exhibit significant error rates, resulting in inaccurate semantic relationships between entities. (ii) global inconsistency: the absence of mechanisms to enforce hierarchical consistency and global coherence during extraction leads to structurally fragmented graphs with poor connectivity. These deficiencies collectively manifest as structural conflicts and semantic ambiguity within the knowledge graph, which subsequently corrupt the retrieval and generation processes. Although recent attempts have been made to refine graph quality via bottom-up clustering-based community summarization (Edge et al., 2024; Gutiérrez et al., 2025; Wang et al., 2025) or topic modeling (Sarthi et al., 2024) to offer a broader, macro-level view of data, these unsupervised methods are vulnerable to error propagation, where inaccuracies in entity relationships are amplified at higher levels of abstraction.
@@ 2.1 Performance Degradation in GraphRAG Systems
GraphRAG models frequently underperform traditional RAG approaches on many real-world tasks. Specifically, our experiments on GraphRAG-Bench (Xiang et al., 2025) demonstrate that while GraphRAG leverages graph structures to enhance recall by retrieving a wider array of potentially relevant passages, this gain is offset by a significant introduction of noisy and ambiguous contextual information. Specifically, GraphRAG methods such as LightRAG and HippoRAG achieve moderate improvements in evidence recall, but exhibit significantly lower context relevance, ranging from 36.86% to 54.61%, compared to Vanilla RAG, which attains 62.87% performance. This suggests that although graph-based retrieval expands the scope of contextual information, it introduces substantial noise that compromises the relevance and reliability of generated answers. For example, in a question-answering task about “climate change impacts”, GraphRAG might retrieve passages related to “economic policies” due to tenuous graph links. In contrast, vanilla RAG preserves tighter alignment with the query context, leading to more accurate and stable outputs.
@@ 2.3 Discussion
The conventional GraphRAG pipeline relies heavily on explicit relation extraction and triple-based knowledge representation. While this approach aims to summarize passages into structured relational forms, it faces two fundamental issues: First, extracting concise and accurate relational triples is computationally expensive and linguistically challenging. Relations expressed in natural language are often complex and context-dependent, and often nuanced or compositional to be accurately distilled into atomic triples; for example, the sentence “Rachel reluctantly agreed to go running with Phoebe” cannot be cleanly reduced to a single atomic triple without losing critical semantic nuance. Second, explicit relation extraction is unnecessary. Aligned entities, rather than relations, serve as the primary anchors connecting information distributed across passages. The original text preserves relational semantics in full context, which can be interpreted dynamically by large language models during inference, without relying on error-prone extraction.
@@ 3.1 Token-free Graph Construction
We construct a hierarchical graph, called Tri-graph, with multiple granularities that is efficient to maintain and update. Given a corpus with a set of passages $\mathcal{P}$}, we first segment each passage into sentences using punctuation (e.g., periods or exclamation marks), obtaining a sentence set $\mathcal{S}$}. We then apply lightweight models (e.g., spaCy (Honnibal et al., 2020)) for named entity recognition (NER) to derive an entity set $\mathcal{E}$}. The passages, sentences, and entities constitute three types of nodes in the graph, denoted as $V_{p}$, $V_{s}$, and $V_{e}$, respectively.
@@ 3.2.1 First Stage: Relevant Entity Activation via Semantic Bridging
Initial entity activation that identifies entities contained in the query $q$ and represents their activation in the knowledge graph as a binary vector:
$\mathbf{a}_{q}=[\mathbf{a}_{q,i}]_{|V_{e}|\times 1},\quad\text{where}\quad\mathbf{a}_{q,i}=\mathbb{1}_{\texttt{sim}(q,e_{i})>0.5}\cdot\texttt{sim}(q,e_{i}).$ (3)
@@
$\mathbf{a}_{q}^{t}=\texttt{MAX}(\sigma_{q}^{T}M\mathbf{a}_{q}^{t-1},\mathbf{a}_{q}^{t-1}),$ (5)
@@
Noteworthy, the vectorized formulation of equation 5 enables $n$-hop entity activation within only $n$ iterations, where $n$ is generally small ($\leq 4$). Each iteration involves two MatMul (matrix multiplication) and one MAX operation, which can be efficiently implemented and executed in parallel on popular machine learning platforms. Furthermore, due to the intrinsic sparsity of the matrix $M$ and the activation vector $\mathbf{a}_{q}^{t}$, we can store it in a sparse format to reduce memory consumption, and reduce computation cost by applying SpMM (sparse matrix multiplication) to replace MatMul.
@@ 4.1 Experimental Setting
Datasets. We first evaluate the effectiveness of LinearRAG on three widely-used multi-hop QA datasets and one domain-specific dataset, including HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (2Wiki) (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), and the Medical dataset from GraphRAG-Bench (Xiang et al., 2025). We follow the same evaluation method as HippoRAG, using the same corpus for retrieval and choosing 1,000 questions from each validation set. This setup allows for a fair comparison between different methods. We also test our approach on the domain-specific Medical dataset from GraphRAG-Bench (Xiang et al., 2025), showing that LinearRAG improves both generation results and retrieval quality.
Baselines. We categorize all the baselines into three groups: (i) Zero-shot LLM Inference: We evaluate several foundational models including LLaMA3 (8B) and LLaMA3 (13B)  (Dubey et al., 2024), as well as GPT-3.5-turbo and GPT-4o-mini (OpenAI, 2023). (ii) We deploy Vanilla RAG methodologies across multiple retrieval configurations (retrieving 1, 3, or 5 top passages), combining semantic-based document retrieval with chain-of-thought reasoning prompts to guide the LLM’s generation process. (iii) State-of-the-art GraphRAG Systems: We compare against leading GraphRAG implementations including KGP  (Wang et al., 2024), G-retriever (He et al., 2024), RAPTOR  (Sarthi et al., 2024), E2GraphRAG (Zhao et al., 2025), LightRAG (Guo et al., 2024), HippoRAG (Gutiérrez et al., 2024), GFM-RAG (Luo et al., 2025), and HippoRAG2 (Gutiérrez et al., 2025). Among these, RAPTOR organizes the corpus into a hierarchical tree graph; G-Retriever, LightRAG, HippoRAG, GFM-RAG, and HippoRAG2 extract triples from passages to build structured graphs; while E2GraphRAG integrates both strategies.
Evaluation Metrics. We evaluate our method using four metrics across two categories. For end-to-end QA performance, following existing work (Wang et al., 2025), we use: 1) Contain-Match Accuracy (Contain-Acc.), which checks if the correct answer appears in the generated response, and 2) GPT-Evaluation Accuracy (GPT-ACC.), an LLM-based metric that assesses whether the predicted answer matches the ground truth. For the Medical dataset, since golden answers consist of lengthy descriptive statements, we only evaluate using GPT-ACC. For retrieval quality assessment, we adopt metrics from GraphRAG-Bench (Xiang et al., 2025): 1) Context Relevance, which measuring semantic alignment between questions and retrieved passages, and 2) Evidence Recall, which evaluating whether the retrieved contents contain all necessary information for the correct answer.
Implementations. For consistency in implementation, all algorithms use the same embedding model, i.e., all-mpnet-base-v2 (Xiao et al., 2023)). We set $k$ = 5 for top $k$ retrieval in all methods. All RAG approaches employ the same LLM (GPT-4o-mini) for both generation and evaluation tasks. All experiments are conducted on the hardware configuration detailed in Appendix C.
@@ Appendix A Datasets
Our experimental evaluation is conducted on four datasets: three established multi-hop benchmark datasets for multi-hop question answering—HotpotQA (Yang et al., 2018), MuSiQue (Trivedi et al., 2022), and 2WikiMultiHopQA (2Wiki) (Ho et al., 2020)—and one domain-specific dataset. Below, we provide a concise overview of each dataset’s key characteristics.
(i) HotpotQA (Yang et al., 2018): A comprehensive benchmark comprising 97k question-answer instances designed to evaluate multi-hop reasoning capabilities. Each question requires models to synthesize information from multiple documents, with up to 2 gold-standard supporting passages provided alongside numerous irrelevant documents. This structure challenges systems to perform effective cross-document inference and evidence selection.
(ii) 2WikiMultiHopQA (2Wiki) (Ho et al., 2020): A multi-hop reasoning benchmark containing 192k questions that necessitate information integration across multiple Wikipedia articles. Each instance requires evidence synthesis from either 2 or 4 specific articles, testing models’ ability to perform structured cross-document reasoning and maintain coherent information flow.
(iii) MuSiQue (Trivedi et al., 2022): A sophisticated multi-hop QA benchmark featuring 25k question-answer pairs that demand 2-4 sequential reasoning steps. Each question requires coherent multi-step logical inference across multiple documents, challenging systems to execute structured reasoning chains while preserving contextual consistency throughout the inference process.
(iv) Medical: A specialized subset derived from GraphRAG-Bench (Xiang et al., 2025), constructed from structured clinical data sourced from the National Comprehensive Cancer Network (NCCN) guidelines. These guidelines provide standardized treatment protocols, drug interaction hierarchies, and diagnostic criteria. The dataset encompasses four tasks of increasing complexity: fact retrieval, complex reasoning, contextual summarization, and creative generation, totaling 4,076 questions across all difficulty levels.

### v4

@@ Abstract
Retrieval-Augmented Generation (RAG) is widely used to mitigate hallucinations of Large Language Models (LLMs) by leveraging external knowledge. While effective for simple queries, traditional RAG systems struggle with large-scale, unstructured corpora where information is fragmented. Recent advances incorporate knowledge graphs to capture relational structures, enabling more comprehensive retrieval for complex, multi-hop reasoning tasks. However, existing graph-based RAG (GraphRAG) methods rely on unstable and costly relation extraction for graph construction, often producing noisy graphs with incorrect or inconsistent relations that degrade retrieval quality. In this paper, we revisit the pipeline of existing GraphRAG systems and propose Linear Graph-based Retrieval-Augmented Generation (LinearRAG), an efficient framework that enables reliable graph construction and precise passage retrieval. Specifically, LinearRAG constructs a relation-free hierarchical graph, termed Tri-Graph, using only lightweight entity extraction and semantic linking, avoiding unstable relation modeling. This new paradigm of graph construction scales linearly with corpus size and incurs no extra token consumption, providing an economical and reliable indexing of the original passages. For retrieval, LinearRAG adopts a two-stage strategy: (i) relevant entity activation via local semantic bridging, followed by (ii) passage retrieval through global importance aggregation. Extensive experiments on four datasets demonstrate that LinearRAG significantly outperforms baseline models. Our code and datasets are available at https://github.com/DEEP-PolyU/LinearRAG.git.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a promising approach to enhance Large Language Models (LLMs) by leveraging external knowledge bases (gao2023retrieval; lewis2020retrieval; zhou-etal-2025-taming; zhang2025faithfulrag). However, existing RAG systems struggle with the complexities of large-scale, unstructured corpora in real-world scenarios, where the relevant information is frequently distributed unevenly across heterogeneous documents. The context retrieved by RAG systems is often voluminous, intricate, and lacks clear organization, leading to issues of variability in accuracy and coherence  (sun2024thinkongraph; zhang2024knowgpt). Although recent advances attempt to manage this by segmenting documents into smaller chunks for efficient indexing (borgeaud2022improving; izacard2023atlas; jiang2023active), this strategy often results in the loss of critical contextual details, impairing retrieval accuracy and reasoning capabilities for complex tasks (han2024retrieval; zhang2025survey).
To address this, Graph Retrieval-Augmented Generation (GraphRAG) (zhang2025survey; procko2024graph; xiao2025graphragbenchchallengingdomainspecificreasoning; xiang2025use) has recently emerged as a powerful paradigm that leverages external structured graphs to model the hierarchical structure of background knowledge (han2024retrieval). Specifically, early work, like RAPTOR (sarthi2024raptor) and Microsoft’s GraphRAG (edge2024local), organizes knowledge through recursive summarization and community detection with LLM-generated synopses, enabling coarse-to-fine retrieval for comprehensive responses. Building on this, recent approaches, including GFM-RAG (luo2025gfm), G-Retriever (he2024g), and LightRAG (guo2024lightrag), integrate specialized encoders and objectives, such as query-dependent GNNs, Prize Collecting Steiner Trees, and dual-level indexing, to improve multi-hop generalization, scalability, and efficiency. More recently, HippoRAG (hipporag) and its enhancement HippoRAG2 (gutiérrez2025hipporag2) draw inspiration from cognitive processes to utilize personalized PageRank for multi-hop retrieval. These strategies significantly improve retrieval precision and contextual depth, enabling LLMs to address complex, multi-hop queries more effectively.
@@
Despite its conceptual promise and theoretical superiority, recent studies reveal that GraphRAG models frequently underperform even naive RAG approaches on many real-world applications (han2025rag; zhou2025depth; xiang2025use; zhuang2025losemb). This performance degradation mainly stems from the poor quality of automatically constructed knowledge graphs. While graph-based retrieval increases recall of relevant knowledge, it concurrently introduces substantial noise and ambiguities into the retrieved contexts, due to errors in graph construction. Specifically, two critical deficiencies undermine graph quality, including (i) local inaccuracy: relation extraction processes exhibit significant error rates, resulting in inaccurate semantic relationships between entities. (ii) global inconsistency: the absence of mechanisms to enforce hierarchical consistency and global coherence during extraction leads to structurally fragmented graphs with poor connectivity. These deficiencies collectively manifest as structural conflicts and semantic ambiguity within the knowledge graph, which subsequently corrupt the retrieval and generation processes. Although recent attempts have been made to refine graph quality via bottom-up clustering-based community summarization (edge2024local; gutiérrez2025hipporag2; wang2025archrag) or topic modeling (sarthi2024raptor) to offer a broader, macro-level view of data, these unsupervised methods are vulnerable to error propagation, where inaccuracies in entity relationships are amplified at higher levels of abstraction.
@@ 2.1 Performance Degradation in GraphRAG Systems
GraphRAG models frequently underperform traditional RAG approaches on many real-world tasks. Specifically, our experiments on GraphRAG-Bench (xiang2025use) demonstrate that while GraphRAG leverages graph structures to enhance recall by retrieving a wider array of potentially relevant passages, this gain is offset by a significant introduction of noisy and ambiguous contextual information. Specifically, GraphRAG methods such as LightRAG and HippoRAG achieve moderate improvements in evidence recall, but exhibit significantly lower context relevance, ranging from 36.86% to 54.61%, compared to Vanilla RAG, which attains 62.87% performance. This suggests that although graph-based retrieval expands the scope of contextual information, it introduces substantial noise that compromises the relevance and reliability of generated answers. For example, in a question-answering task about “climate change impacts”, GraphRAG might retrieve passages related to “economic policies” due to tenuous graph links. In contrast, vanilla RAG preserves tighter alignment with the query context, leading to more accurate and stable outputs.
@@ 2.3 Discussion
wThe conventional GraphRAG pipeline relies heavily on explicit relation extraction and triple-based knowledge representation. While this approach aims to summarize passages into structured relational forms, it faces two fundamental issues: First, extracting concise and accurate relational triples is computationally expensive and linguistically challenging. Relations expressed in natural language are often complex and context-dependent, and often nuanced or compositional to be accurately distilled into atomic triples; for example, the sentence “Rachel reluctantly agreed to go running with Phoebe” cannot be cleanly reduced to a single atomic triple without losing critical semantic nuance. Second, explicit relation extraction is unnecessary. Aligned entities, rather than relations, serve as the primary anchors connecting information distributed across passages. The original text preserves relational semantics in full context, which can be interpreted dynamically by large language models during inference, without relying on error-prone extraction.
@@ 3.1 Token-free Graph Construction
We construct a hierarchical graph, called Tri-graph, with multiple granularities that is efficient to maintain and update. Given a corpus with a set of passages $\mathcal{P}$}, we first segment each passage into sentences using punctuation (e.g., periods or exclamation marks), obtaining a sentence set $\mathcal{S}$}. We then apply lightweight models (e.g., spaCy (honnibal2020spacy)) for named entity recognition (NER) to derive an entity set $\mathcal{E}$}. The passages, sentences, and entities constitute three types of nodes in the graph, denoted as $V_{p}$, $V_{s}$, and $V_{e}$, respectively.
@@ 3.2.1 First Stage: Relevant Entity Activation via Semantic Bridging
Initial entity activation that identifies entities contained in the query $q$ and represents their activation in the knowledge graph as a sparse vector. Specifically, we first use SpaCy to extract entities from $q$, denoted as $E_{q}$. Then for each extracted entity, we find the most similar entity in the knowledge graph. The initial activation score of the matched entity is set to its similarity score:
$\mathbf{a}_{q}=[\mathbf{a}_{q,i}]_{|V_{e}|\times 1},\quad\text{where}\quad\mathbf{a}_{q,i}=\mathbb{1}_{i=\arg\max_{e_{j}\in V_{e}}\texttt{sim}(e_{q},e_{j})}\cdot\texttt{sim}(e_{q},e_{i}).$ (3)
@@
$\mathbf{a}_{q}^{t}=\texttt{MAX}(M^{T}(\sigma_{q}\odot(M\mathbf{a}_{q}^{t-1})),\mathbf{a}_{q}^{t-1}),$ (5)
@@
Noteworthy, the vectorized formulation of equation 5 enables $n$-hop entity activation within only $n$ iterations, where $n$ is generally small ($\leq 4$). Each iteration involves three MatMul (matrix multiplication) and one MAX operation, which can be efficiently implemented and executed in parallel on popular machine learning platforms. Furthermore, due to the intrinsic sparsity of the matrix $M$ and the activation vector $\mathbf{a}_{q}^{t}$, we can store it in a sparse format to reduce memory consumption, and reduce computation cost by applying SpMM (sparse matrix multiplication) to replace MatMul.
@@ 4.1 Experimental Setting
Datasets. We first evaluate the effectiveness of LinearRAG on three widely-used multi-hop QA datasets and one domain-specific dataset, including HotpotQA (yang2018hotpotqa), 2WikiMultiHopQA (2Wiki) (2wikimqa), MuSiQue (trivedi2022MuSiQue), and the Medical dataset from GraphRAG-Bench (xiang2025use). We follow the same evaluation method as HippoRAG, using the same corpus for retrieval and choosing 1,000 questions from each validation set. This setup allows for a fair comparison between different methods. We also test our approach on the domain-specific Medical dataset from GraphRAG-Bench (xiang2025use), showing that LinearRAG improves both generation results and retrieval quality.
Baselines. We categorize all the baselines into three groups: (i) Zero-shot LLM Inference: We evaluate several foundational models including LLaMA3 (8B) and LLaMA3 (13B)  (dubey2024llama3), as well as GPT-3.5-turbo and GPT-4o-mini (openai2023gpt4). (ii) We deploy Vanilla RAG methodologies across multiple retrieval configurations (retrieving 1, 3, or 5 top passages), combining semantic-based document retrieval with chain-of-thought reasoning prompts to guide the LLM’s generation process. (iii) State-of-the-art GraphRAG Systems: We compare against leading GraphRAG implementations including KGP  (wang2024knowledge), G-retriever (he2024g), RAPTOR  (sarthi2024raptor), E2GraphRAG (zhao20252GraphRAG), LightRAG (guo2024lightrag), HippoRAG (hipporag), GFM-RAG (luo2025gfm), and HippoRAG2 (gutiérrez2025hipporag2). Among these, RAPTOR organizes the corpus into a hierarchical tree graph; G-Retriever, LightRAG, HippoRAG, GFM-RAG, and HippoRAG2 extract triples from passages to build structured graphs; while E2GraphRAG integrates both strategies.
Evaluation Metrics. We evaluate our method using four metrics across two categories. For end-to-end QA performance, following existing work (Wang et al., 2025), we use: 1) Contain-Match Accuracy (Contain-Acc.), which checks if the correct answer appears in the generated response, and 2) GPT-Evaluation Accuracy (GPT-ACC.), an LLM-based metric that assesses whether the predicted answer matches the ground truth. For the Medical dataset, since golden answers consist of lengthy descriptive statements, we only evaluate using GPT-ACC. For retrieval quality assessment, we adopt metrics from GraphRAG-Bench (xiang2025use): 1) Context Relevance, which measuring semantic alignment between questions and retrieved passages, and 2) Evidence Recall, which evaluating whether the retrieved contents contain all necessary information for the correct answer.
Implementations. For consistency in implementation, all algorithms use the same embedding model, i.e., all-mpnet-base-v2 (bge_embedding)). We set $k$ = 5 for top $k$ retrieval in all methods. All RAG approaches employ the same LLM (GPT-4o-mini) for both generation and evaluation tasks. All experiments are conducted on the hardware configuration detailed in Appendix C.
@@ Appendix A Datasets
Our experimental evaluation is conducted on four datasets: three established multi-hop benchmark datasets for multi-hop question answering—HotpotQA (yang2018hotpotqa), MuSiQue (trivedi2022MuSiQue), and 2WikiMultiHopQA (2Wiki) (2wikimqa)—and one domain-specific dataset. Below, we provide a concise overview of each dataset’s key characteristics.
(i) HotpotQA (yang2018hotpotqa): A comprehensive benchmark comprising 97k question-answer instances designed to evaluate multi-hop reasoning capabilities. Each question requires models to synthesize information from multiple documents, with up to 2 gold-standard supporting passages provided alongside numerous irrelevant documents. This structure challenges systems to perform effective cross-document inference and evidence selection.
(ii) 2WikiMultiHopQA (2Wiki) (2wikimqa): A multi-hop reasoning benchmark containing 192k questions that necessitate information integration across multiple Wikipedia articles. Each instance requires evidence synthesis from either 2 or 4 specific articles, testing models’ ability to perform structured cross-document reasoning and maintain coherent information flow.
(iii) MuSiQue (trivedi2022MuSiQue): A sophisticated multi-hop QA benchmark featuring 25k question-answer pairs that demand 2-4 sequential reasoning steps. Each question requires coherent multi-step logical inference across multiple documents, challenging systems to execute structured reasoning chains while preserving contextual consistency throughout the inference process.
(iv) Medical: A specialized subset derived from GraphRAG-Bench (xiang2025use), constructed from structured clinical data sourced from the National Comprehensive Cancer Network (NCCN) guidelines. These guidelines provide standardized treatment protocols, drug interaction hierarchies, and diagnostic criteria. The dataset encompasses four tasks of increasing complexity: fact retrieval, complex reasoning, contextual summarization, and creative generation, totaling 4,076 questions across all difficulty levels.